OpenStudy (sasogeek):

if \(\large (-2)^0=1 \) what is \(\large -2^0 \) ?

OpenStudy (anonymous):

-2^0 means -1 * 2^0 which as you know is -1 * (1) = -1

OpenStudy (jiteshmeghwal9):

then it's will also 1

OpenStudy (jiteshmeghwal9):

\((-1)^0.(2)^0\) 1.1=1

OpenStudy (anonymous):

remember when doing -x^y it means take x to the y and minus it, not take (-x) to the y

OpenStudy (jiteshmeghwal9):

\(-2^0\)=1

OpenStudy (anonymous):

(-2)^0 = 1

OpenStudy (anonymous):

not -2^0

OpenStudy (shubhamsrg):

@jiteshmeghwal9 (-2)^0 and -2^0 ,,both are diff..

OpenStudy (jiteshmeghwal9):

\(- \times\)\(2^0\) =-1

OpenStudy (jiteshmeghwal9):

now it is right @apple_pi

OpenStudy (anonymous):

yes :D